# Sensor-based Multi-Robot Search and Coverage with Spatial Separation in Unstructured Environments

### 04 Mar 2024 v1

@@ Abstract
Multi-robot systems have increasingly become instrumental in tackling search and coverage problems. However, the challenge of optimizing task efficiency without compromising task success still persists, particularly in expansive, unstructured environments with dense obstacles. This paper presents an innovative, decentralized Voronoi-based approach for search and coverage to reactively navigate these complexities while maintaining safety. This approach leverages the active sensing capabilities of multi-robot systems to supplement GIS (Geographic Information System), offering a more comprehensive and real-time understanding of the environment. Based on point cloud data, which is inherently non-convex and unstructured, this method efficiently generates collision-free Voronoi regions using only local sensing information through spatial decomposition and spherical mirroring techniques. Then, deadlock-aware guided map integrated with a gradient-optimized, centroid Voronoi-based coverage control policy, is constructed to improve efficiency by avoiding exhaustive searches and local sensing pitfalls. The effectiveness of our algorithm has been validated through extensive numerical simulations in high-fidelity environments, demonstrating significant improvements in both task success rate, coverage ratio, and task execution time compared with others.
@@ I Introduction
Multi-robot systems are extensively employed in search and coverage tasks across diverse applications, including gas leak detection, pollution source identification, and search and rescue operations [1, 2]. These systems can efficiently locate Targets of Interest (ToI) in unknown and unstructured environments through interactive data collection. Building on their capabilities, integration with Geographic Information Systems (GIS) offers new opportunities for improving efficiency and accuracy of such operations [3].
Figure 1: Utilizing GIS-generated heat maps as prior information for multi-robot systems in both post-disaster search and rescue and area monitoring for disaster prevention.
@@
On the other hand, some reactive approaches compute search policies that focus on the immediate rewards associated with ToI, thereby enhancing efficiency through the local sensing range. These methods, rigorously examined in the context of Centroidal Voronoi Tessellation (CVT) in multi-robot coverage problems [2, 5, 6, 7] offer numerous advantages, including asynchronous operation, distributed execution, adaptability, and verifiable asymptotic correctness [8]. However, they are susceptible to local minima due to short-sighted decisions, which can lead to the failure to complete the task. Furthermore, the majority of existing methods are confined to two-dimensional, well-structured environments [9]. They fall short in accounting for unstructured settings, including raw sensor data, where robots are often trapped in the presence of non-convex obstacles when attempting to expedite their greedy search [10, 11, 12].
In this paper, we introduce a novel decentralized Voronoi-based active search and coverage strategy for unknown and unstructured environments with guaranteed safety. This research is driven by scenarios where robots aim to balance coverage efficiency and success rate while adhering to safety constraints. We aim to develop a cooperative control policy that enables individual robots to make intelligent decisions while enhancing overall team search efficiency. First, we utilize realistic point cloud data as input, characterized by its highly non-convex and unstructured nature. By integrating spatial decomposition and spherical mirroring techniques, collision-free Voronoi regions are efficiently generated using only local sensing and position information from other robots. Second, Explored information is employed to construct a deadlock-aware guided map to refine subsequent processes for optimal decision-making. This map is integrated with centroid Voronoi based coverage control policy to improve efficiency by avoiding exhaustive searches and local sensing pitfalls. Third, The effectiveness of our proposed method has been extensively validated through numerical simulations in high-fidelity large-scale environments, including cluttered, thin-structural, and narrow settings with guaranteed safety. Compared with other Voronoi-based methods, the approach has significantly improved successful coverage ratio and time.
@@ II Related Work
This work on search and coverage integrates contributions related to its functionality, building upon collision avoidance methods and our previous research on coverage control [12].
A crucial consideration in collision avoidance is the formulation of obstacles derived from point cloud data. Most existing methods rely on using high resources to maintain a cost map, such as the Euclidean Distance Transform (EDT) map, to calculate the minimum distance to obstacles for optimization [13, 14, 15], which may suffer from time-consuming computations and redundant processes [16]. Alternatively, some strategies depend on down-sampling and spatial partitioning using convex clusters, [17]. Unfortunately, such techniques may sacrifice the free space or lead to unsafe areas [18]. Instead, our approach directly generates safe Voronoi cells on a voxel-based environment representation, inspired by the concept presented in [19]. The advantages lie in threefold: 1) It operates exclusively on point cloud data, eliminating the need for a post-processed map and ensuring the preservation of raw sensor measurement fidelity; 2) It does not assume the convexity or structure of the environment, making it particularly suitable for real-world sensors influenced by noise and adaptable for a range of scenarios; 3) The strategy requires only the sensing of other robots’ positions, which simplifies communication requirements and bolsters system robustness.
@@ IV Sensor-based Coverage Control with Deadlocks and Collision Avoidance
Figure 2: Instance of multi-robot search and coverage mission in unstructured environments with three sparsely distributed information sources.
@@ IV-A2 Robot-obstacle Collision Avoidance Hyperplane
Moreover, due to the monotonically decreasing nature of the function $\|F(\cdot)\|$, points that are closer to the robot are transformed further away. Since there is no point outside the convex hull $\Omega_{c}$, it implies that the area inside the corresponding origin points $\mathbf{q}_{1},\dots,\mathbf{q}_{m_{o}}$, is obstacle-free. With the described spherical mirroring operation, the extraction of points that are necessary for generating robot-obstacle collision avoidance hyperplanes becomes a straightforward task for an unordered point cloud.
@@ IV-B Safe Region Construction
Based on above, $\mathbf{a}_{io}$ can be quickly calculated by solving the following low-dimensional quadratic programming problem:
@@
We then shift the hyper-plane to be tight with the target. Thus, $b_{io}=\min{\mathbf{a}_{io}^{\mathrm{T}}\mathbf{q}_{o}}$.
@@ V Simulations
This section analyzes the applicability of the proposed method across diverse settings, including real forests and indoor office datasets, and evaluates the performance of computational time and collision distance. Comparisons between the proposed method and state-of-the-art baselines will be introduced. The simulations are conducted on a laptop equipped with an Intel i7-9700 CPU and 16GB of RAM. A team of robots is simulated through multiprocessing. For real-world scenario validation, we utilized the Robot Operating System (ROS) with C++ and Python programming languages. MATLAB is employed for comparative analysis. Velocity constraints are established with a maximum limit of $u_{\text{max}}=2.5$m/s. The time interval for each replanning is $\Delta t=0.1$s. The obstacles are inflated by a robot radius, $r_{i}$= 0.25m. The grid map resolution is also set to 0.25m for all axes. The sensor’s radius is restricted to a range of 15m. If $\|\mathbf{u}\|<0.01$m/s, the robot is considered to have converged.
@@ V-B Comparisons of Coverage Performance
We undertake comparative evaluations with other Voronoi-based coverage methods: 1) OAVC method in [11], formulates secure areas by computing tangential boundary lines to circular obstacles and adheres to the gradient of target information; 2) Adaptive coverage method in [9], utilizes a repulsive density function concerning detected obstacles, ensuring that the newly computed centroid is positioned away from the obstacle. To be fair, the obstacles are modeled as the same shape in [11] We evaluate these methods over 30 trials, each with randomly initialized starting positions within an 80m $\times$ 80m $\times$ 16m workspace. The ToI is defined by a Gaussian distribution featuring three peaks. The performance metrics are as follows: 1) Coverage ratio: The ratio of detected peaks to the total number of peaks; 2) Success rate: The fraction of trials where robots successfully converge to the ToI without experiencing deadlock or collisions within a limited time (30s) or until all peaks are detected; 3) Average time: Task execution time, considering only successful trials. The results are shown in Tab. I.
@@ VI Conclusions
In this work, the proposed coverage method, which incorporates collision avoidance without entrapment, is effectively applicable across a range of unstructured settings. The approach leverages spatial decomposition and sphere flipping transformation to construct a safe Voronoi region in complex environments with considerable point cloud data. The NF is integrated to steer the robot away from pitfalls while simultaneously directing it towards the ToI. Comparisons with others highlight that our approach performs extremely well, with high success ratios and guaranteed safety. Our method also extends the traditional coverage control method, move-to-centroid policy, into more realistic scenarios. Simulations demonstrate that our approach is capable of generating more robust, efficient, and secure solutions for multi-robot coverage problems.

### 10 Apr 2024 v2

# Sensor-based Multi-Robot Coverage Control with Spatial Separation in Unstructured Environments

@@ Abstract
Multi-robot systems have increasingly become instrumental in tackling coverage problems. However, the challenge of optimizing task efficiency without compromising task success still persists, particularly in expansive, unstructured scenarios with dense obstacles. This paper presents an innovative, decentralized Voronoi-based coverage control approach to reactively navigate these complexities while guaranteeing safety. This approach leverages the active sensing capabilities of multi-robot systems to supplement GIS (Geographic Information System), offering a more comprehensive and real-time understanding of environments like post-disaster. Based on point cloud data, which is inherently non-convex and unstructured, this method efficiently generates collision-free Voronoi regions using only local sensing information through spatial decomposition and spherical mirroring techniques. Then, deadlock-aware guided map integrated with a gradient-optimized, centroid Voronoi-based coverage control policy, is constructed to improve efficiency by avoiding exhaustive searches and local sensing pitfalls. The effectiveness of our algorithm has been validated through extensive numerical simulations in high-fidelity environments, demonstrating significant improvements in task success rate, coverage ratio, and task execution time compared with others.
@@ I Introduction
Multi-robot systems are extensively employed in coverage tasks across diverse applications, including gas leak detection, pollution source identification, and search and rescue operations [1, 2]. These systems can efficiently locate Targets of Interest (ToI) in unknown and unstructured environments through interactive data collection. Building on their capabilities, integration with Geographic Information Systems (GIS) offers new opportunities for improving efficiency and accuracy of such operations [3].
Figure 1: GIS provides maps as prior information for multi-robot systems in post-disaster search and rescue.
@@
On the other hand, some reactive approaches compute search policies that focus on the immediate rewards associated with ToI, thereby enhancing efficiency through the local sensing range. These methods, rigorously examined in the context of Centroidal Voronoi Tessellation (CVT) in multi-robot coverage problems [2, 5, 6, 7] offer numerous advantages, including asynchronous operation, distributed execution, adaptability, and verifiable asymptotic correctness [8]. However, they are susceptible to local minima due to short-sighted decisions, which can lead to the failure to complete the task. Furthermore, the majority of existing methods are confined to two-dimensional, well-structured environments [9]. They fall short in accounting for unstructured settings, including raw sensor data, where robots are often trapped in the presence of non-convex obstacles when attempting to expedite their greedy coverage policy [10, 11, 12].
In this paper, we introduce a novel decentralized Voronoi-based active coverage policy for unknown and unstructured environments with guaranteed safety. This research is driven by scenarios where robots aim to balance coverage efficiency and success rate while adhering to safety constraints. We aim to develop a cooperative control policy that enables individual robots to make intelligent decisions while enhancing overall team efficiency. Firstly, we utilize realistic point cloud data as input, characterized by its highly non-convex and unstructured nature. By integrating spatial decomposition and spherical mirroring techniques, collision-free Voronoi regions are efficiently generated using only local sensing and position information from other robots. Secondly, explored information is employed to construct a deadlock-aware guided map to refine subsequent processes for optimal decision-making. This map is integrated with centroid Voronoi based coverage control policy to improve efficiency by avoiding exhaustive searches and local sensing pitfalls. Thirdly, the effectiveness of our proposed method has been extensively validated through numerical simulations in high-fidelity large-scale environments, including cluttered, thin-structural, and narrow settings. Compared with other Voronoi-based methods, the approach has significantly improved the successful coverage ratio and time.
@@ II Related Work
This work integrates contributions related to its functionality, building upon collision avoidance methods and our previous research on coverage control [12].
A crucial consideration in collision avoidance is the formulation of obstacles derived from point cloud data. Most existing methods rely on using high resources to maintain a cost map, such as the Euclidean Distance Transform (EDT) map, to calculate the minimum distance to obstacles for optimization [13, 14, 15], which may suffer from time-consuming computations and redundant processes [16]. Alternatively, some strategies depend on down-sampling and spatial partitioning using convex clusters, [17]. Unfortunately, such techniques may sacrifice the free space or lead to unsafe areas [18]. Instead, our approach directly generates safe Voronoi cells on a voxel-based environment representation, inspired by the concept presented in [19, 20]. The advantages lie in threefold: 1) It operates exclusively on point cloud data, eliminating the need for a post-processed map and ensuring the preservation of raw sensor measurement fidelity; 2) It does not assume the convexity or structure of the environment, making it particularly suitable for real-world sensors influenced by noise and adaptable to a range of scenarios; 3) The strategy requires only the sensing of other robots’ positions, which simplifies communication requirements and bolsters system robustness.
@@ IV Sensor-based Coverage Control with Deadlocks and Collision Avoidance
Figure 2: Multi-robot coverage mission in unstructured environments with three sparsely distributed information sources.
@@ IV-A2 Robot-obstacle Collision Avoidance Hyperplane
Moreover, due to the monotonically decreasing nature of the function $\|F(\cdot)\|$, points that are closer to the robot are transformed further away. Since there is no point outside the convex hull $\Omega_{c}$, it implies that the area inside the corresponding origin points $\mathbf{q}_{1},\dots,\mathbf{q}_{m_{c}}$, is obstacle-free. With the described spherical mirroring operation, the extraction of points that are necessary for generating robot-obstacle collision avoidance hyperplanes becomes a straightforward task for an unordered point cloud.
@@ IV-B Safe Region Construction
Based on above, $\mathbf{a}_{io}$ can be quickly calculated by solving the following low-dimensional quadratic programming (QP) problem:
@@
We then shift the hyper-plane to be tight with the target. Thus, $b_{io}=\min{\mathbf{a}_{io}^{\mathrm{T}}\mathbf{q}_{o}}$. Solving a QP problem requires $O(N^{3})$ time, where $N$ denotes the number of decision variables [26]. These decision variables are linearly correlated with the number of obstacle points $m_{c}$ in our method. Consequently, the total computational complexity for each robot, influenced by the local density of obstacles, is $O(m_{c}^{3})$. Leveraging advanced optimization utilities, such as CVXOPT, we can solve the problem in tens of milliseconds.
@@ V Simulations
This section analyzes the applicability of the proposed method across diverse and unstructured settings, including real forests and indoor office datasets, and evaluates the performance of computational time and collision distance. Besides, benchmarking our method against established coverage control protocols can help quantify the improvements in task efficiency. The accompanying video provides a clear visual aid to understand our method’s mechanics and advantages. [^1] https://youtu.be/wOunvjGHhBQ The simulations are conducted on a laptop equipped with an Intel i7-9700 CPU and 16GB of RAM. A team of robots is simulated through multiprocessing. For real-world scenario validation, We utilized the Robot Operating System (ROS) with C++ and Python programming languages. MATLAB is employed for comparative analysis. Velocity constraints are established with a maximum limit of $u_{\text{max}}=2.5$m/s. The time interval for each replanning is $\Delta t=0.1$s. The obstacles are inflated by a robot radius, $r_{i}$= 0.25m. The grid map resolution is also set to 0.25m for all axes. The sensor’s radius is restricted to a range of 15m. If $\|\mathbf{u}\|<0.01$m/s, the robot is considered to have converged.
@@ V-B Comparisons of Coverage Performance
We undertake comparative evaluations with other Voronoi-based coverage methods: 1) OAVC method in [11], formulates secure areas by computing tangential boundary lines to circular obstacles and adheres to the gradient of target information; 2) Adaptive coverage method in [9], utilizes a repulsive density function concerning detected obstacles, ensuring that the newly computed centroid is positioned away from the obstacle. To be fair, the obstacles are modeled as the same shape in [11]. We evaluate these methods over 30 trials, each with randomly initialized starting positions within an 80m $\times$ 80m $\times$ 16m workspace. The ToI is defined by a Gaussian distribution featuring three peaks. The performance metrics are as follows: 1) Coverage ratio: The ratio of detected peaks to the total number of peaks; 2) Success rate: The fraction of trials where robots successfully converge to the ToI without experiencing deadlock or collisions within a limited time (30s) or until all peaks are detected; 3) Average time: Task execution time, considering only successful trials. The results are shown in Tab. I.
@@ VI Conclusions
In this work, our proposed method, which incorporates collision avoidance without entrapment, is effectively applicable across a range of unstructured settings. The approach leverages spatial decomposition and sphere flipping transformation to construct a safe Voronoi region in complex environments with considerable point cloud data. The NF is integrated to steer the robot away from pitfalls while simultaneously directing it towards the ToI. Comparisons with others highlight that our approach performs extremely well, with high success ratios and guaranteed safety. Our method also extends the traditional coverage control method, move-to-centroid policy, into more realistic scenarios. Future work includes testing our algorithm on physical robot swarms to address hardware-specific challenges and to further validate its real-world applicability.